# UMBCLU at SemEval-2024 Task 1A and 1C: Semantic Textual Relatedness with and without machine translation

### 20 Feb 2024 v1

@@ Abstract
This paper describes the system we developed for SemEval-2024 Task 1, “Semantic Textual Relatedness for African and Asian Languages.” The aim of the task is to build a model that can identify semantic textual relatedness (STR) between two sentences of a target language belonging to a collection of African and Asian languages. We participated in Subtasks A and C and explored supervised and cross-lingual training leveraging large language models (LLMs). Pre-trained large language models have been extensively used for machine translation and semantic similarity. Using a combination of machine translation and sentence embedding LLMs, we developed a unified STR model, TranSem, for subtask A and fine-tuned the T5 family of models on the STR data, FineSem, for use in subtask C. Our model results for 7 languages in subtask A were better than the official baseline for 3 languages and on par with the baseline for the remaining 4 languages. Our model results for the 12 languages in subtask C resulted in $1^{st}$ place for Africaans, $2^{nd}$ place for Indonesian, and $3^{rd}$ place for English with low performance for the remaining 9 languages.
UMBCLU at SemEval-2024 Task 1A and 1C: Semantic Textual Relatedness with and without machine translation
@@ 1 Introduction
The objective of the SemEval 2024 Task 1 is to build and evaluate models capable of identifying relatedness between sentence pairs from a language. Both sentences in a sentence pair are from a single language taken from a selection of 14 languages belonging to 5 language groups. The task is divided into 3 sub-tasks, where each sub-task is aimed at studying a different type of model training: supervised (Subtask A), unsupervised (Subtask B), and cross-lingual (Subtask C). The languages and the data made available for each subtask are different. Extensive details about the languages, language groups, data, and the collection process can be found in the task description paper Ousidhoum et al. (2024a). A detailed description of the task, sub-tasks, and datasets can be found in the shared task description Ousidhoum et al. (2024b).
@@
Semantic relatedness helps with understanding language meaning (Jarmasz and Szpakowicz, 2012; Miller, 1992) and is useful in many areas of natural language processing such as word-sense disambiguation (Banerjee and Pedersen, 2003), machine translation (Bracken et al., 2017), sentence representation (Reimers et al., 2019), question answering and summarization (Wang et al., 2022) and memory association (Antony et al., 2022; Osgood, 1949). However, semantic relatedness has been mostly restricted to finding word relatedness (Feng et al., 2017; Budanitsky and Hirst, 2006), leading to a lack of sentence relatedness datasets. At the sentence level, relatedness has been limited to similarity, leading to a set of semantic textual similarity (STS) benchmark datasets and models. Similarity is a subset of relatedness and provides a very restricted view of STR (Abdalla et al., 2021). The current shared task aims to develop models and resources for sentence-level relatedness in several languages (Ousidhoum et al., 2024b).
In this paper, we describe our models for the supervised and cross-lingual training subtasks A & C. For subtask A, our aim was to develop a single unified model trained on all the training data for all languages and find the STR of a sentence pair from any language. In such a model, we also could find STR for pairs where each sentence is from a different language. To train the unified model, we use machine translation to convert sentences from all the languages into English. Our target language is English because of the available sentence embedding models in English. For subtask C, our goal was to fine-tune a pre-trained LLM on the STS benchmark dataset and then the English STR dataset and evaluate its usefulness in the cross-domain setting. T5 was a natural choice as the STS dataset is part of the model’s pretraining data.
Our primary contributions in this paper are as follows:
  1. 1.
    We explore if a single unified model can be trained based on translations of the source language.
  2. 2.
    We explore the use of translation models for augmenting data.
  3. 3.
    We explore whether a T5 trained on the English STR dataset works well for any of the target languages in the cross-domain setting.
Our code is publicly available at https://github.com/dipta007/SemEval24-Task8
@@ 2 System Overview
After exploring models and datasets available in languages we understand [^1] Of the 14 languages, the authors are proficient in English, Hindi & Telugu languages., we realized the dearth of resources in these languages. Our goal was to explore how to leverage resources available in English for the languages in the shared task. The result of our effort is two different systems for the supervised and cross-lingual training sub-tracks A & C. Our goal for subtask A was to develop a single model for all languages and explore the effectiveness of machine translation. Our goal for subtask C was to explore the effectiveness of transfer learning from English STS to other languages.
2.1 Model for Subtask A
In this section, we describe our translation-based sentence embedding system to use in the supervised setting. Our goal is to translate the sentences from the various languages to English. We train a single model on the translated data and perform inference on the translated data. Assuming the translated data reflects the semantic meaning of the source language accurately, the derived relatedness value from our model reflects the STR between a sentence pair in the source language. The performance of our model allows us to evaluate the machine translation performance by comparing it with the performance of the English dataset, which does not undergo translation.
2.1.1 Translation to English & Data Augmentation
While a number of SOTA translation models are available, we selected Meta’s “No Language Left Behind (NLLB) open-source models” that provide direct high-quality translations for 200 languages where many are low resource languages(Costa-jussà et al., 2022). We used four of the translation models222 facebook/nllb-200-3.3B, facebook/nllb-200-1.3B, facebook/nllb-200-distilled-1.3B, facebook/nllb-200-distilled-600M ranging from 600 million to 3.3 billion parameters for translation to English. We translated the training data for all languages in Subtask A, except Amharic and Algerian Arabic, with each of the models. Thus, we obtained 4 translated datasets for each language, leading to a 4-fold augmentation of our data. We excluded Amharic and Algerian Arabic from our study as none of these models supported the two languages. Aiming to obtain the best translation features at inference time, for the test data, we translated using only the largest model (facebook/all-200-3.3 B).
2.1.2 Our Model (TranSem)
Figure 1: Overview of our TranSem model architecture (Inspired by Reimers et al. (2019)). The encoder ($\theta$) is shared, and the diamond box represents the loss function. The input to the cosine similarity loss is the encoded sentence pairs ($x_{1},x_{2}$) and the label ($y$).
@@ 2.1.2 Our Model (TranSem)
We experimented with several sentence encoding models and chose DistilRoberta333sentence-transformers/all-distilroberta-v1 to submit results for the competition leaderboard. This is a distilled version of RoBERTa (Liu et al., 2019) fine-tuned on sentence-level datasets and suitable for clustering and semantic searches, which we further fine-tuned on our translated and augmented dataset. We selected this model for submitting the results based on our primary validation (details on § 4.1.1, § 4.1.2, and § 4.1.3). After experimenting with different pooling mechanisms of mean, max, and CLS tokens, we found that mean pooling works well for our setting. This aligns with earlier findings, which showed that mean pooling produced encodings that lead to better performance on downstream tasks.
2.2 Our Model for Subtask C (FineSem)
In this section, we describe our use of the T5 family of models for cross-lingual training. T5 is a text-to-text transformer model that uses transfer learning where the model is pre-trained on a mixture of unsupervised and supervised tasks in a multi-task setting Raffel et al. (2020). One of the supervised tasks used in the T5 model training was the semantic textual similarity benchmark (STS-B) dataset trained as a regression classification problem. We fine-tuned the pre-trained T5-base, T5-large, and XL models ranging from (660 million to 3 billion parameters) on the STR English training data, treating using the same format as the STS-B dataset and using regression classification. Since this model is not allowed for use on the English test set for subtask C (according to the rules for the subtask), we fine-tuned another pretrained-T5 model on the Spanish STR training data for use on the English test set. We also used these fine-tuned T5 models for Subtask A and compared them with our TranSem model.
@@ 3.1 HyperParameters
To train our model, we have used AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 1e-5 and weight decay of 0.01. We use an effective batch size of 32 (batch size 16 with 2 steps of gradient accumulation) (for TransSem) and a batch size of 16 (for FineSem). We train our model infinitely with an early stopping patience of 10 on the validation Spearman Correlation score.
@@ 3.2 Infrastructure
All experiments are conducted on an NVIDIA Quadro RTX 8000 with 48GB of VRAM and A100 80GB. We utilize the PyTorch Lightning library  [^4] https://lightning.ai/ for conducting the experiments and Weight & Biases  [^5] https://wandb.ai/ for logging purposes (for the TransSem Model) and HuggingFace Transformers (for the FineSem Model).
@@ 4 Results
In this section, first, we report our results and analysis on Subtask A (§ 4.1), and then we report the same on Subtask C (§ 4.2).
4.1 Subtask A
We use the official baseline (described on Ousidhoum et al. (2024b)). They have used LaBSE (Feng et al., 2020) fine-tuned on the provided training dataset. Throughout the paper, we refer to this model as $baseline$.
We have fine-tuned our model (shown in Fig. 1) on the training dataset. Throughout the paper, we refer to this model as $TranSem$.
@@ 4.1.1 Effect of Batch Size
We tried various batch sizes (results in Table 1) and showed that our batch size selections are fairly good (32 for task A and 8 for SubTask C).
@@ 4.1.2 Effect of Encoder Pooling
For the official results we submitted, we used mean pooling as it showed good performance in most languages. In Table 2, we compare results with 3 different pooling operations.
@@
4.1.3 Effect of Different Models
We provide constrastive results with several other models in Table 3. For the sentence embedding models, we reported official results using distilroberta-v1. We also submitted some comparative results with a T5-XL model fine-tuned on the STR training datasets without any translation (last row in the Table 2)
@@ 4.1.3 Effect of Different Models
fine-tuned T5-XL .8381 .6056 .6713 .1827 .7714 .6089 .6818 .6228
Table 3: Model Performance (Spearman Correlation Coefficient) on Subtask A test set. $TranSem$ shows results submitted before the official deadline, $baseline$ shows official baseline results, and the rest are contrastive results for various models we tried. The bold value represents the best score within a specific section, and the underlined value signifies the best score across all sections. The last row “fine-tuned T5-XL” shows the results of a T5-XL model fine-tuned on the STR training data for each language
4.2 Subtask C
We use the English fine-tuned models for inference on all language test sets except English and use the Spanish fine-tuned models for inference on the English test set. We submitted official results using the FineSem model, a T5-XL model fine-tuned for 3 epochs. Our results in Table 4 show the performance of fine-tuned T5-base, Large, and XL models fine-tuned for 10 epochs. We report the best performance for each language, along with the checkpoint saved at a particular epoch number that gave the best performance.
C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11
Model Afr Arq Amh Eng Hau Hin Ind Kin Arb Ary Punj
Baseline .6200 .4600 .8400 .8000 .6200 .7600 .4700 .5700 .6100 .4000 -.050
FineSem .8223 1263 .0430 .7875 .4569 .1552 .5153 .4836 .0354 -1.038 -.078
T5-B .7926 .1799 .2037 .7448 .4784 .2472 .4517 .4437 .0307 .0307 -.308
T5-L .8068 -.035 .0965 .7794 .5052 .2089 .4846 .5529 .0539 .0331 -.012
Table 4: Model performance (Spearman Correlation Coefficient) on Subtask C test sets. All language test sets (except English) use the FineSem models trained on the English training set. The English test set uses the FineSem models trained on the Spanish training set. T5-B, T5-L point to fine-tuning other model sizes of T5 for reporting constrastive results.
@@ 5 Related Work
In this section, we will present the previous research conducted in the fields of Machine Translation (see § 5.1), Sentence Embedding (see § 5.2) and Semantic Similarity (see § 5.3).
@@ 6 Conclusion & Future Work
We developed two different models and showed how the models performed in supervised and cross-domain training tasks in 14 languages. We explored the use of machine translation, sentence encoders, and SST-B style training with T5 models. Our models improved over the official baseline for some of the languages. For computational purposes, we have excluded using more recent models like mistral-7b (Jiang et al., 2023), which have outperformed most of the open-source and close-source models in various benchmarks (Zheng et al., 2024). Also, we believe that future work can use the prompt-based LLMs  [^6] https://chat.openai.com/ for translation.
@@ 7 Limitations
We acknowledge our work has the following limitations. We use several pre-trained large language models in our experiments. It is well known that these models can echo biases and misinformation either implicitly or explicitly. We did not control for any of these when training them on the STR datasets. In addition, the STR datasets may also echo several biases related to social groups, cultural groups, race, gender, behavioral, and perceptual differences of annotators. We did not explore or control for any of these biases in our work. As a result, our work carries the limitations of both the models and the datasets we used.

### 12 Apr 2024 v2

# UMBCLU at SemEval-2024 Task 1: Semantic Textual Relatedness with and without machine translation

@@ Abstract
The aim of SemEval-2024 Task 1, “Semantic Textual Relatedness for African and Asian Languages” is to develop models for identifying semantic textual relatedness (STR) between two sentences using multiple languages (14 African and Asian languages) and settings (supervised, unsupervised, and cross-lingual). Large language models (LLMs) have shown impressive performance on several natural language understanding tasks such as multilingual machine translation (MMT), semantic similarity (STS), and encoding sentence embeddings. Using a combination of LLMs that perform well on these tasks, we developed two STR models, TranSem and FineSem, for the supervised and cross-lingual settings. We explore the effectiveness of several training methods and the usefulness of machine translation. We find that direct fine-tuning on the task is comparable to using sentence embeddings and translating to English leads to better performance for some languages. In the supervised setting, our model performance is better than the official baseline for 3 languages with the remaining 4 performing on par. In the cross-lingual setting, our model performance is better than the baseline for 3 languages (leading to $1^{st}$ place for Africaans and $2^{nd}$ place for Indonesian), is on par for 2 languages and performs poorly on the remaining 7 languages.
UMBCLU at SemEval-2024 Task 1: Semantic Textual Relatedness with and without machine translation
@@ 1 Introduction
The objective of the SemEval 2024 Task 1 is to build and evaluate models capable of identifying relatedness between a sentence pair. Sentence pairs from 14 African and Asian languages belonging to 5 language groups are annotated for relatedness and released for model development. The task is divided into 3 tracks targeting different types of model training: supervised (Track A), unsupervised (Track B), and cross-lingual (Track C). Each track targets a different subset of languages. Extensive details about the languages, language groups, and the data collection process are provided in the task description paper Ousidhoum et al. (2024a). A detailed description of the shared task, tracks, and datasets are provided in the shared task description paper Ousidhoum et al. (2024b).
@@
Semantic relatedness helps with understanding language meaning (Jarmasz and Szpakowicz, 2012; Miller, 1995; Antony et al., 2022; Osgood, 1949) and is useful in many areas of natural language processing such as word-sense disambiguation (Banerjee and Pedersen, 2003), machine translation (Bracken et al., 2017) and sentence representation (Reimers et al., 2019; Wang et al., 2022) which have numerous applications. Until recently, semantic relatedness has been mostly restricted to finding word relatedness (Feng et al., 2017; Budanitsky and Hirst, 2006), leading to a lack of sentence-relatedness datasets. At the sentence level, relatedness has been limited to similarity, providing a restricted view of STR (Abdalla et al., 2021). The current shared task aims to broaden the scope of sentence relatedness and extend it to several languages with the goal of encouraging model and resource development in these languages (Ousidhoum et al., 2024b).
Recent advancements in multi-lingual translation and the availability of models for obtaining high-quality sentence embeddings allowed us to explore the effectiveness of machine translated data. Using various sentence embedding models to encode data translated into English, we trained a model, TranSem, to find the relatedness score between sentence pairs. Although the task requires the sentence pair to be from the same language, our model can handle sentences from two different languages. Our second model, FineSem directly fine-tunes a T5 model (T5 is already fine-tuned on the STS benchmark) on the STR task using both untranslated and translated data to explore the usefulness of translation. We use both these models to evaluate languages in Track A. For Track C languages, we use a T5 model fine-tuned only on the english STR data. For evaluating the English dataset in the cross-lingual track, we use a T5 model fine-tuned on the Spanish dataset.
Our contributions to the STR task are as follows: We 1) develop unified models for STR to work with all languages. 2) participate in supervised and cross-lingual tracks. 3) explore the usefulness of machine translation. 4) explore data augmentation using machine translation. Our code is publicly available at https://github.com/dipta007/SemEval24-Task8
@@ 2 System Overview
After exploring models and datasets available in the languages we understand [^1] Of the 14 languages, the authors are proficient in English, Hindi & Telugu languages., we realized the dearth of resources available in these languages. To leverage resources available in English, we translated the 13 non-English languages into English. Assuming the translated data accurately reflects the semantic meaning of the source language, the derived relatedness value from our model for a translated sentence pair should reflect the STR between the sentence pair in the source language. This section describes our machine translation process and models, TranSem and FineSem. Besides using different training strategies, these models can use both translated and untranslated data.
2.1 Translation to English & Data Augmentation
We use Meta’s “No Language Left Behind (NLLB) open-source models” that provide direct high-quality translations for 200 languages with many low-resource languages (Costa-jussà et al., 2022). We use four of the translation models222 facebook/nllb-200-3.3B, facebook/nllb-200-1.3B, facebook/nllb-200-distilled-1.3B, facebook/nllb-200-distilled-600M ranging from 600 million to 3.3 billion parameters. Using each of the 4 models, we translated the training data for all languages in track A, except Amharic and Algerian Arabic, and obtained 4 translated datasets for each language. None of the 4 models we used supported Amharic, Algerian Arabic or Punjabi. We decided against translating track C languages with 3 of 12 unsupported languages. Using 4 different model translations gave us a 4-fold augmentation of the training data. We translated the test data using only the largest model (facebook/all-200-3.3 B) to obtain the best translation features.
2.2 TranSem Model
Figure 1: Overview of TranSem model architecture (Inspired by Reimers et al. (2019)). The encoder ($\theta$) is shared, and the diamond box represents the loss function. The encoded sentence pairs ($x_{1},x_{2}$) and the label ($y$) are the input to the cosine similarity loss.
@@ 2.2 TranSem Model
We experimented with several sentence encoding models for encoding our translated and augmented training dataset. We chose DistilRoberta333sentence-transformers/all-distilroberta-v1 to submit results for the competition leaderboard based on our primary validation (details on § 4.1.1, § 4.1.2, and § 4.1.3). This is a distilled version of RoBERTa (Liu et al., 2019) fine-tuned on sentence-level datasets and suitable for clustering and semantic searches, which we further fine-tuned on our translated and augmented dataset. The sentence-t5-xl embedding model was chosen to compare the effectiveness of sentence embeddings as opposed to the direct fine-tuning used in the FineSem model. After experimenting with different pooling mechanisms of mean, max, and CLS tokens, we found that mean pooling works well for our setting. This aligns with earlier findings, which show that mean pooling produced encodings lead to better performance on downstream tasks.
2.3 FineSem Model
T5 Raffel et al. (2020) is a transformer model that uses transfer learning; the model trained on “Colossal Clean Crawled Corpus” is fine-tuned on a mixture of 8 downstream unsupervised and supervised tasks converting them into a unified text-to-text task setting. The T5 model is available in several sizes, of which we use the base, large, and XL models ranging from (660 million to 3 billion parameters). One of the supervised tasks used in the T5 model training is the semantic textual similarity benchmark (STS-B) dataset trained as a regression classification problem. We use the STS task setting to train on the track A STR training datasets using 3 different options: separate T5 models trained on individual languages, a single model trained on all 14 languages (without translation), and a single model trained on the translated and augmented dataset (for 12 languages). These settings allow us to contrast the effectiveness of direct fine-tuning with the sentence embedding-based transem model and the usefulness of machine translation.
From the T5 models fine-tuned on the individual languages, we use the English and Spanish fine-tuned models (we refer to these models as the English and Spanish models) for evaluating the cross-lingual Track C languages. We use the English model to evaluate development and test data from all languages except English and the Spanish model to evaluate the English data.
@@ 3.1 HyperParameters
We train our models using AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 1e-5 and weight decay of 0.01. We use an effective batch size of 32 (batch size 16 with 2 steps of gradient accumulation) (for TransSem) and a batch size of 16 (for FineSem).
We train our transem model infinitely with an early stopping patience of 10 on the validation Spearman Correlation score. We train the finesem model for 10 epochs (2 epochs for the model trained on the translated and augmented model) and checkpoint the models at the end of every epoch. We evaluate the dev sets for each language against these 10 checkpoints. We evaluate the corresponding test data using the checkpoint which provides the best performance on the dev data for a language.
@@ 3.2 Infrastructure
All experiments were conducted on an NVIDIA Quadro RTX 8000 with 48GB of VRAM and A100 80GB. We utilize the PyTorch Lightning library  [^4] https://lightning.ai/ for conducting the experiments and Weight & Biases  [^5] https://wandb.ai/ for logging purposes (for the TransSem Model) and HuggingFace Transformers (for the FineSem Model).
@@ 4 Results
We first report our results and analysis on Track A languages (§ 4.1), and then on Track C (§ 4.2). We use the official baseline (Ousidhoum et al. (2024b)) that used LaBSE (Feng et al., 2020) fine-tuned on the provided training dataset and refer to this model as $baseline$.
4.1 Track A Languages
This section discusses our findings using various model settings with the TranSem model.
@@ 4.1.1 Effect of Batch Size
Comparing performance with various batch sizes (results are shown in Table 1), we show that our batch size selections are fairly good (32 for task TranSem and 8 for FineSem).
@@ 4.1.2 Effect of Encoder Pooling
In Table 2, we compare performance using 3 different pooling operations. We used mean pooling for the official results we submitted, as it showed good performance in most languages.
@@
4.1.3 Effect of Sentence Embedding Models
In Table 3, we provide contrastive results with several sentence embedding models used in TranSem. For official results, we submitted results from the distilroberta-v1 sentence embedding model (results for some of the languages are from the FineSem model fine-tuned on individual STR training datasets).
@@ 4.1.3 Effect of Sentence Embedding Models
FineSem-Individual .8385 .6335 .7175 .2211 .7647 .6900 .6085 .6391
FineSem-Unified .8438 .6369 .6837 .3878 .6265 .7040 .6993 .6546
FineSem-Translated .8105 .6383 .7133 .8608 .7403 .6663 .8152 .7493
Table 3: Model Performance (Spearman Correlation Coefficient) on Subtask A test set. $TranSem$ shows results submitted before the official deadline, $baseline$ shows official baseline results, and the rest are contrastive results for our various models. The best scores within each section are bolded, and best scores across all sections are underlined.
4.1.4 Usefulness of Machine Translation and Direct Fine-tuning
We compare the performance of the FineSem models fine-tuned using the 3 data options (results are shown in Table 3). FineSem-Individual shows the performance of T5-XL models fine-tuned on the individual datasets. Unified and Translated models show the performance of the two T5-XL models fine-tuned on the untranslated data and translated$+$augmented data. The model trained on untranslated data performs poorly on the Marathi dataset, but performs on par with the other models indicating that we may not need to translate all languages to English. We find that direct fine-tuning with the translated and augmented data is comparable with the TranSem model using various sentence embeddings.
4.2 Track C Languages
In Table 4, we compare the performance of various T5 models on the track C languages. We submitted official results using our T5-XL based FineSem model (FineSem-LB) where the results are obtained using the checkpoint after the third epoch. With the same model we also report with the approach where we use the checkpoint which results in the best performance on the development data for a given language. These results are shown as FineSem-XL. We compare these results with the T5-base and T5-large based FineSem models. We bold the best scores for easy readability but underline scores that are better than the baseline. Among our models the overall performance of the XL model is better and this model improves upon the baseline for Afrikaans, Indonesian and Spanish.
C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11 C12
Model Afr Arq Amh Eng Hau Hin Ind Kin Arb Ary Pan Esp avg
Baseline .7900 .4600 .8400 .8000 .6200 .7600 .4700 .5700 .6100 .4000 -.050 .6200 .5742
FineSem-LB .8223 1263 .0430 .7875 .4569 .1552 .5153 .4836 .0354 -.0375 -.0775 .6089 .3266
FineSem-XL .8164 1023 .0373 .7889 .4561 .1594 .5279 .4128 .$0$0219 -.0817 .6259 .3246
FineSem-L .8007 -.0515 .0112 .7752 .4831 .1764 .4419 .5094 .0154 .0331 -.0591 .6605 .3164
FineSem-B .7802 .1799 .2543 .7448 .4784 .2404 .4517 .3861 .0527 .0268 -.0520 .6289 .3477
Table 4: Model performance (Spearman Correlation Coefficient) on Track C test sets. All language test sets (except English) use the FineSem models trained on the English training set. The English test set uses the FineSem models trained on the Spanish training set. The best scores among our models are bolded. Scores better than baseline are Underlined.
@@ 5 Related Work
In this section, we present previous research conducted in the fields of Machine Translation (see § 5.1), Sentence Embedding (see § 5.2) and Semantic Similarity (see § 5.3).
@@ 6 Conclusion & Future Work
We developed two different models and showed how the models performed in supervised and cross-domain training tasks in 14 languages. We explored using machine translation, sentence encoders, and SST-B style training with T5 models. Our models improved over the official baseline for some of the languages. For computational purposes, we have excluded using more recent models like mistral-7b (Jiang et al., 2023), which have outperformed most of the open-source and close-source models in various benchmarks (Zheng et al., 2024). For future work, we intend to explore prompting for STR and prompt-based LLMs  [^6] https://chat.openai.com/ for translation.
7 Disclaimer
We did not use AI assistants to write any part of our paper or code. All writing is original and produced by the authors.
@@ 8 Limitations
We acknowledge our work has the following limitations. We use several pre-trained LLMs in our experiments. It is well known that these models can echo biases and misinformation either implicitly or explicitly. We did not control for any of these when training them on the STR datasets. In addition, the STR datasets may also echo several biases related to social groups, cultural groups, race, gender, behavioral, and perceptual differences of annotators. We did not explore or control for any of these biases in our work. As a result, our work carries the limitations of both the models and the datasets we used.